# VideoLLaMA 3: Frontier Multimodal Foundation Models for Image and Video Understanding

### v3

@@ 4.2.2 Benchmarks
We assess VideoLLaMA3’s general video understanding capabilities through established benchmarks: (1) Multi-Choice Video Question Answering (MC-VQA) tasks, including MVBench (li2023mvbench, ), VideoMME (fu2024video, ), EgoSchema (mangalam2024egoschema, ), and Perception-Test (patraucean2024perception, ). (2) Open-Ended Video Question Answering (OE-VQA) tasks, including ActivityNet-QA (yu2019activitynet, ) and VCGBench (Maaz2023VideoChatGPT, ). This evaluation suite follows the protocol of VideoLLaMA2 (damonlpsg2024videollama2, ). We also run evaluations on MMVU (zhao2025mmvu, ) which includes both the task types mentioned above.
@@
To further examine the capacity of VideoLLaMA3 to process and comprehend long-form video content, we assess performance on three long-video understanding (LVU) benchmarks: (1) MLVU (MLVU, ): diverse long-video understanding tasks for videos ranging from 3 minutes to more than 2 hours, (2) LongVideoBench (wu2024longvideobench, ): video reasoning over the referred context within long video-language interleaved inputs, and (3) LVBench (wang2024lvbench, ): extreme long video understanding.
@@
To assess the temporal awareness and reasoning capabilities of VideoLLaMA3, we conduct evaluations on the following tasks: (1) Temporal Perception and Reasoning tasks, including TempCompass (liu2024tempcompass, ) and NextQA (xiao2021next, ); and (2) Temporal Sentence Grounding task on Charades-STA (gao2017tall, ) benchmark, with mean Intersection over Union (mIoU) metric.
@@ 4.4 Ablation Study
 clip-vit-large-patch14-336 radford2021learning 61.50 56.28 18.32 24.86 1668.41
dfn5B-clip-vit-h-14-378 fang2023data 62.70 56.87 16.40 23.09 1665.35
siglip-so400m-patch14-384 Zhai2023SigmoidLF 62.92 57.12 22.44 31.32 1667.92
In MLLMs, the embeddings of pre-trained vision encoder should be trained to align with embeddings of LLMs. Therefore, the representation performance of vision encoder is crucial to the final performance of MLLMs. In this work, we study the impact of different vision encoders. Specifically, we compare three pre-trained transformer-based vision encoders: CLIP radford2021learning , DFN fang2023data , and SigLIP Zhai2023SigmoidLF . Due to the computation limitation, we perform the study on the subset of the whole dataset. Also, to investigate the performance of the original pre-trained weights, we fix the weights of vision encoders and keep the visual inputs as the fixed resolution, which is the same as the pretrained resolution of the vision encoder ($336\times 336$ for CLIP, $378\times 378$ for DFN, and $384\times 384$ for SigLIP,). The training has three stages: 1) Training projector with LLaVA-Pretrain-558K (liu2023improvedllava, ); 2) Tuning all parameters with our recaptioned COYO data; 3) SFT with LLaVA-SFT-665K (li2024llava, ). The comparison results are shown in Table. 9. SigLIP outperforms the other two vision encoders, especially in fine-grained understanding tasks involving texts. Based on this ablation study, we choose the pretrained SigLIP as our base vision encoder, and then adapt it to taking dynamic resolutions as inputs.
@@ 5 Related Work
Multimodal LLMs for Native Video Understanding. Early video MLLMs primarily relied on sparsely sampled frames and simple connectors, such as MLPs (maaz2023video, ; lin2023video, ; ataallah2024minigpt4, ), discrete visual tokenizers (jin2024video, ), and Q-formers (zhang2023video, ; li2024mvbench, ), to link visual encoders with large language models. Subsequent models propose various methods to overcome token limitations and support long-form video understanding. For example,  (zhang2024longva, ) directly extends the context window of LLMs to achieve long video understanding, while others (shen2024longvu, ; xu2024pllava, ; jin2024chat, ; zhang2024video, ; xu2024slowfast, ; weng2025longvlm, ; song2024moviellm, ; li2025llama, ; tan2024koala, ) introduce video token compression techniques that perform pooling across spatial, temporal, or both dimensions. While most approaches utilize image-based encoders (zohar2024apollo, ; song2024moviechat+, ; chen2024sharegpt4video, ; fei2024video, ; wang2024tarsierrecipestrainingevaluating, ; hong2024cogvlm2, ; liu2024kangaroo, ; liu2024oryx, ; chen2024timemarker, ), some incorporate video-specific encoders to better capture temporal dependencies (wang2024internvideo2, ; jung2024pegasus, ; Maaz2024VideoGPT+, ; cheng2024videollama, ).
More recent works (sun2024video, ; cheng2024videollama, ; ye2025cat, ; shu2023audio, ; su2023pandagpt, ; lyu2023macaw, ; panagopoulou2023x, ; han2024onellm, ; yu2024crema, ; tang2024avicuna, ) extend beyond visual inputs by incorporating audio, using separate encoders for each modality and integrating them through an LLM decoder. These models leverage joint instruction tuning on video-audio datasets (Yang2022AVQAAD, ; AlAmri2019AudioVS, ; ye2024cat, ; leng2024curse, ) to capture interactions between visual and auditory information. Additionally, recent advances in streaming video understanding focus on real-time processing (zhang2024flash, ; qian2024streaming, ; Chen2024VideoLLMonlineOV, ; Zhang2024InternLMXComposer25OmniLiveAC, ; Liu2024StreamChatCW, ), employing techniques like adaptive memory and incremental processing for tasks such as live event detection and real-time captioning. Previous works shen2024longvu ; zhang2024video ; tan2024koala ; zohar2024apollo ; wang2024internvideo2 ; cheng2024videollama typically follow a training recipe that involves an alignment phase, followed by supervised fine-tuning, with instruction-tuning datasets (zhang2024video, ; Maaz2023VideoChatGPT, ; li2023mvbench, ; liu2024oryx, ; chen2024sharegpt4video, ) often being video dominant. However, we propose a vision-centric training paradigm to enhance video understanding capabilities by focusing on large-scale image understanding pre-training. This approach leverages high-quality image-text datasets to build robust vision encoders that are then adapted for video tasks.
@@
Previous studies jiang2024mantisinterleavedmultiimageinstruction ; li2024llavanextinterleavetacklingmultiimagevideo ; li2024llavaonevision ; mplug3 have demonstrated that general MLLMs with robust image understanding capabilities can achieve remarkable performance on video understanding tasks, even with minimal or no dedicated video training data. These findings highlight the effectiveness of task transfer from images to videos, showcasing the models’ strong video comprehension and cross-scenario adaptability.
Furthermore, Qwen2-VL wang2024qwen2 adopts a unified framework for processing both images and videos, enhancing the model’s visual perception capabilities. Models such as Qwen2-VL, InternVL-2 chen2023internvl , and InternVL-2.5 internvl2.5 , which scale both model sizes (ranging from 1 billion to 78 billion parameters) and the volume of training data, have achieved highly competitive performance in both image and video understanding tasks. To address the challenges of processing longer video inputs, recent studies longllava ; internlmxcomposer25 have proposed solutions such as adapting model architectures by incorporating a hybrid design of Mamba and Transformer blocks or training with extensive long video datasets to support extended input and output sequences.
Recent studies vita ; vita15 ; llama3 have integrated text, image, and video modalities with audio and speech modalities to improve models’ video understanding and cross-scenario performance. Additionally, Aria aria utilizes a fine-grained mixture-of-experts decoder, which enables more efficient training and inference compared to dense decoders when handling multimodal inputs.

### v4

@@ 4.2.2 Benchmarks
We assess VideoLLaMA3’s general video understanding capabilities through established benchmarks: (1) Multi-Choice Video Question Answering (MC-VQA) tasks, including MVBench [27], VideoMME [127], EgoSchema [128], and Perception-Test [129]. (2) Open-Ended Video Question Answering (OE-VQA) tasks, including ActivityNet-QA [130] and VCGBench [26]. This evaluation suite follows the protocol of VideoLLaMA2 [46]. We also run evaluations on MMVU [131] which includes both the task types mentioned above.
@@
To further examine the capacity of VideoLLaMA3 to process and comprehend long-form video content, we assess performance on three long-video understanding (LVU) benchmarks: (1) MLVU [132]: diverse long-video understanding tasks for videos ranging from 3 minutes to more than 2 hours, (2) LongVideoBench [133]: video reasoning over the referred context within long video-language interleaved inputs, and (3) LVBench [134]: extreme long video understanding.
@@
To assess the temporal awareness and reasoning capabilities of VideoLLaMA3, we conduct evaluations on the following tasks: (1) Temporal Perception and Reasoning tasks, including TempCompass [135] and NextQA [136]; and (2) Temporal Sentence Grounding task on Charades-STA [108] benchmark, with mean Intersection over Union (mIoU) metric.
@@ 4.4 Ablation Study
 clip-vit-large-patch14-336 [137] 61.50 56.28 18.32 24.86 1668.41
dfn5B-clip-vit-h-14-378 [138] 62.70 56.87 16.40 23.09 1665.35
siglip-so400m-patch14-384 [54] 62.92 57.12 22.44 31.32 1667.92
In MLLMs, the embeddings of pre-trained vision encoder should be trained to align with embeddings of LLMs. Therefore, the representation performance of vision encoder is crucial to the final performance of MLLMs. In this work, we study the impact of different vision encoders. Specifically, we compare three pre-trained transformer-based vision encoders: CLIP [137], DFN [138], and SigLIP [54]. Due to the computation limitation, we perform the study on the subset of the whole dataset. Also, to investigate the performance of the original pre-trained weights, we fix the weights of vision encoders and keep the visual inputs as the fixed resolution, which is the same as the pretrained resolution of the vision encoder ($336\times 336$ for CLIP, $378\times 378$ for DFN, and $384\times 384$ for SigLIP,). The training has three stages: 1) Training projector with LLaVA-Pretrain-558K [55]; 2) Tuning all parameters with our recaptioned COYO data; 3) SFT with LLaVA-SFT-665K [38]. The comparison results are shown in Table. 9. SigLIP outperforms the other two vision encoders, especially in fine-grained understanding tasks involving texts. Based on this ablation study, we choose the pretrained SigLIP as our base vision encoder, and then adapt it to taking dynamic resolutions as inputs.
@@ 5 Related Work
Multimodal LLMs for Native Video Understanding. Early video MLLMs primarily relied on sparsely sampled frames and simple connectors, such as MLPs [139, 12, 13], discrete visual tokenizers [140], and Q-formers [141, 142], to link visual encoders with large language models. Subsequent models propose various methods to overcome token limitations and support long-form video understanding. For example,  [14] directly extends the context window of LLMs to achieve long video understanding, while others [143, 144, 145, 25, 146, 147, 148, 149, 150] introduce video token compression techniques that perform pooling across spatial, temporal, or both dimensions. While most approaches utilize image-based encoders [15, 16, 28, 17, 151, 152, 18, 19, 153, 154], some incorporate video-specific encoders to better capture temporal dependencies [20, 21, 22, 23].
More recent works [24, 23, 155, 156, 157, 158, 159, 160, 161, 162] extend beyond visual inputs by incorporating audio, using separate encoders for each modality and integrating them through an LLM decoder. These models leverage joint instruction tuning on video-audio datasets [163, 164, 165, 166] to capture interactions between visual and auditory information. Additionally, recent advances in streaming video understanding focus on real-time processing [167, 168, 169, 170, 171], employing techniques like adaptive memory and incremental processing for tasks such as live event detection and real-time captioning. Previous works [143, 25, 150, 15, 20, 23] typically follow a training recipe that involves an alignment phase, followed by supervised fine-tuning, with instruction-tuning datasets [25, 26, 27, 19, 28] often being video dominant. However, we propose a vision-centric training paradigm to enhance video understanding capabilities by focusing on large-scale image understanding pre-training. This approach leverages high-quality image-text datasets to build robust vision encoders that are then adapted for video tasks.
@@
Previous studies [93, 172, 29, 30] have demonstrated that general MLLMs with robust image understanding capabilities can achieve remarkable performance on video understanding tasks, even with minimal or no dedicated video training data. These findings highlight the effectiveness of task transfer from images to videos, showcasing the models’ strong video comprehension and cross-scenario adaptability.
Furthermore, Qwen2-VL [7] adopts a unified framework for processing both images and videos, enhancing the model’s visual perception capabilities. Models such as Qwen2-VL, InternVL-2 [31], and InternVL-2.5 [34], which scale both model sizes (ranging from 1 billion to 78 billion parameters) and the volume of training data, have achieved highly competitive performance in both image and video understanding tasks. To address the challenges of processing longer video inputs, recent studies [32, 33] have proposed solutions such as adapting model architectures by incorporating a hybrid design of Mamba and Transformer blocks or training with extensive long video datasets to support extended input and output sequences.
Recent studies [173, 174, 4] have integrated text, image, and video modalities with audio and speech modalities to improve models’ video understanding and cross-scenario performance. Additionally, Aria [8] utilizes a fine-grained mixture-of-experts decoder, which enables more efficient training and inference compared to dense decoders when handling multimodal inputs.